# PAC-Bayesian Adversarially Robust Generalization Bounds for Graph Neural Network

### v1

@@ Abstract
Graph neural networks (GNNs) have gained popularity for various graph-related tasks. However, similar to deep neural networks, GNNs are also vulnerable to adversarial attacks. Empirical studies have shown that adversarially robust generalization has a pivotal role in establishing effective defense algorithms against adversarial attacks. In this paper, we contribute by providing adversarially robust generalization bounds for two kinds of popular GNNs, graph convolutional network (GCN) and message passing graph neural network, using the PAC-Bayesian framework. Our result reveals that spectral norm of the diffusion matrix on the graph and spectral norm of the weights as well as the perturbation factor govern the robust generalization bounds of both models. Our bounds are nontrivial generalizations of the results developed in (Liao et al., 2020) from the standard setting to adversarial setting while avoiding exponential dependence of the maximum node degree. As corollaries, we derive better PAC-Bayesian robust generalization bounds for GCN in the standard setting, which improve the bounds in (Liao et al., 2020) by avoiding exponential dependence on the maximum node degree.
statistical learning theory, robust generalization bounds, PAC-Bayesian analysis
@@ 1 Introduction
Graph neural network (GNN) (Scarselli et al., 2009) has developed rapidly in recent years, due to the excellent performance in a variety of applications with the graph-structure data. Generally, GNNs are neural networks that combine local aggregation of node features and non-linear transformations to predict the node-level labels or graph-level labels, in semi-supervised learning (Kipf & Welling, 2017) and supervised learning (Pan et al., 2015, 2016; Zhang et al., 2018; Ying et al., 2018), respectively. With the rapid development of GNN, there is an increasing interest in understanding the generalization (Scarselli et al., 2018; Verma & Zhang, 2019; Du et al., 2019; Garg et al., 2020; Liao et al., 2020; Esser et al., 2021; Ju et al., 2023) and other theoretical properties of GNN.
Szegedy et al. (2013) and Goodfellow et al. (2015) reveal that deep neural network (DNN) can be easily fooled by adversarial input samples and that adversarial training can improve the generalization of DNN. GNN also has a similar neural network structure and can be vulnerable to adversarial attacks on graph data. Previous work focuses on how to attack on GNN (Dai et al., 2018; Zügner et al., 2018; Bojchevski & Günnemann, 2019; Zügner et al., 2020; Liu et al., 2019) and how to defend GNN against adversarial attacks (Zhang & Zitnik, 2020; Wu et al., 2019; Entezari et al., 2020; Zhu et al., 2019; Tang et al., 2020), but the theoretical analysis is relatively limited.
A recent line of work discusses the generalization of GNN in the standard (non-adversarial) setting, and derive certain bounds on complexity notations including Vapnik–Chervonenkis (VC) dimension and Rademacher complexity that are used to characterize the generalization error, i.e., the disparity between the performance on training and test datasets of a model. Specifically, Scarselli et al. (2018) derive an upper bound of order $\mathcal{O}(p^{4}n^{2})$ for the VC dimension of GNNs with sigmoid activations, where $p$ is the number of parameters and $n$ is the number of nodes. Verma & Zhang (2019) derive a generalization bound for single-layer graph convolutional neural network (GCN) by analyzing the algorithmic stability of stochastic gradient descent, where the bound depends on the number of algorithm iterations and the largest absolute eigenvalue of the graph convolution filter. Du et al. (2019) consider an infinitely wide multi-layer GNN trained by gradient descent based on neural tangent kernels (Jacot et al., 2018), and derive a generalization bound for this particular model. These two results are constrained to specific settings and do not apply to other scenarios. Garg et al. (2020) present the pioneering data-dependent generalization bounds for GNNs employing Rademacher complexity, with a scaling factor of $\mathcal{O}\left(hd^{l}\sqrt{l\log\left(hd\right)}\right)$. Here, $d$ is the maximum degree of the graph, while $h$ and $l$ denote the width and depth of the network, respectively.
@@
In seeking to grasp the constrained robust generalization abilities of GNNs in the adversarial setting, in this paper, we provide adversarially robust generalization bounds for two kinds of popular GNNs, GCN (Kipf & Welling, 2017) and MPGNN (Dai et al., 2016; Jin et al., 2019), using PAC-Bayesian framework. Our result reveals that spectral norm of the diffusion matrix on the graph, spectral and Frobenius norm of the weights, as well as the perturbation factor govern the robust generalization bounds of both models. Our bounds for GCN complement the findings in the standard setting for DNN in (Neyshabur et al., 2017), as well as the results for GCN in (Liao et al., 2020). All the bounds of these studies have the same spectrally-normalized term of the weights and the dependencies on both depth and width:
@@
To our knowledge, our derived PAC-Bayesian adversarially robust generalization bounds may be the first kind for GNN in the adversarial robustness settings. Even in the standard setting, our bounds improve the bounds in (Liao et al., 2020) from $\mathcal{O}\left(ld^{\frac{l-1}{2}}\sqrt{h\log(lh)}\right)$ to $\mathcal{O}(l\sqrt{h\log(lh)})$ for GCN, and $\mathcal{O}\left(ld^{l-2}\sqrt{h\log(lh)}\right)$ to $\mathcal{O}(l\left\|P_{G}\right\|_{2}^{l-2}\sqrt{h\log(lh)})$ for MPGNN, avoiding the exponential dependence on the maximum degree $d$ of the graph. Meanwhile, our results remove the extra first-order and second-order Lipschitz-continuous assumptions in (Ju et al., 2023) for MPGNN, exhibiting comparable tightness on the dependency on the depth, width and maximum degree. In the context of adversarial robustness, the derived generalization bounds incorporate an additional factor $\epsilon$ that represents the intensity of the perturbation and is in general unavoidable.
@@
The rest of the paper are organized as follows. In Section 2, we briefly review some other related works. In Section 3, we present the problem setup and the main results. We outline the proof techniques in Section 4. Finally, we discuss the limitations and some open problems in Section 5.
@@ 2 Related Work
Generalization Bounds. Classical results rely on complexity notions of neural networks to regulate generalization bounds, employing measures such as VC-dimension as seen in (Anthony & Bartlett, 2009), or Rademacher complexity as discussed in (Bartlett & Mendelson, 2002). Norm-margin-based generalization bounds are developed using Rademacher complexity for fully-connected neural network in (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018). Neyshabur et al. (2015) establish a generalization bound with an exponential dependence $2^{l}$ on the network depth $l$ using Frobenius norms of parameters. Bartlett et al. (2017) showed a margin based generalization bound that depends on spectral norm and $1$-norm of the layers of the networks, reducing the dependence on depth. Golowich et al. (2018) provide bounds on their Rademacher complexity assuming norm constraints on the parameter matrix of each layer, refining the dependency on depth to $\sqrt{l}$. Neyshabur et al. (2017) establish a PAC-Bayesian framework to bound the generalization gap for fully-connected neural network, where the bound depends on the product of the spectral norm of the weights in each layer as well as the Frobenius norm of the weights. As mentioned above, the generalization bounds on GNN are derived using different tools such as VC-dimension (Scarselli et al., 2018), stability analysis (Verma & Zhang, 2019), Rademacher complexity (Garg et al., 2020) and PAC-Bayesian analysis (Liao et al., 2020; Ju et al., 2023).
@@
Adversarial Attack and Defense. Zügner et al. (2018) focus on the adversarial attacks against GCN and provide an algorithm to generate adversarial perturbations targeting on the node features and the graph structures. Dai et al. (2018) propose a reinforcement learning based attack method for adversarial attacks against GNN by modifying the structures of the graph. Zügner et al. (2020) develop a general attack on the training procedure of GCN using meta-gradients. Bojchevski & Günnemann (2019) propose a principled strategy for adversarial attacks on unsupervised node embeddings. There are also some researches focusing on the defense methods. Wu et al. (2019) provide a defense method by pre-processing the adjacency matrix of the graph. The methods in (Zhu et al., 2019; Entezari et al., 2020) defend against attacks by absorbing or dropping the noise information. Tang et al. (2020) use transfer learning to improve the robustness of GNN. Zhang & Zitnik (2020) develop a general algorithm to defend against a variety of training-time attacks that perturb the graph structure.
@@
Adversarial Generalization in DNN. A line of work presents the adversarially robust generalization bounds on DNNs using Rademacher complexity (Khim & Loh, 2018; Yin et al., 2019; Awasthi et al., 2020; Gao & Wang, 2021; Xiao et al., 2022). Khim & Loh (2018) derive robust generalization bounds for linear and neural network via Rademacher complexity of a transformed function. Yin et al. (2019) established generalization bounds for both linear models and one-hidden-layer neural networks based on the surrogate loss introduced in (Raghunathan et al., 2018) considering vector-$\infty$-norm-additive perturbation. Awasthi et al. (2020) give upper and lower bounds for Rademacher complexity of linear hypotheses with vector-$r$-norm-additive perturbation for $r\geq 1$, and also provide upper bounds on Rademacher complexity of one hidden layer neural networks, directly applying to the original network. Gao & Wang (2021) and Farnia et al. (2019) show robust generalization bounds for DNNs, considering specific attacks. Xiao et al. (2022) and Mustafa et al. (2022) employ two different techniques, respectively, to calculate the covering number that can be used to estimate the Rademacher complexity, for adversarial function classes of DNNs. Further, Xiao et al. (2023) present a robust generalization bound on DNN using PAC-Bayesian framework.
3 Adversarially Robust Generalization Bounds
@@
In this section, we first introduce the problem setup for multi-class graph classification problem with GCN and MPGNN, and preliminary results on PAC-Bayesian analysis. We then present our main results: improved generalization bounds in standard settings and adversarially robust generalization bounds for GCN and MPGNN. Finally, we compare our results with existing results.
3.1 Preliminaries
3.1.1 Problem Setup
@@
3.1.2 Background of PAC-Bayesian Analysis
@@
Neyshabur et al. (2017) provide the lemma to give a margin-based generalization bound derived from the PAC-Bayesian bound. Liao et al. (2020) improve the lemma and present the lemma in a two-side form. Both of the above lemmas are proposed in the standard settings. We present Lemma 3.2 of the two-side form with the robust margin loss in the adversarial settings. We provide the proof sketch of Lemma 3.2 in Section 4 while the proof details of Theorems 3.1 and 3.2 will be given in Sections A.2.1 and A.2.2, respectively.
3.2 Main Results
@@
3.2.1 Generalization Bounds for GCN
We now present the improved generalization bounds for GCN in standard setting.
@@ 3.2.1 Generalization Bounds for GCN
3.2.2 Generalization Bounds for MPGNN

### v2

@@ Abstract
Graph neural networks (GNNs) have gained popularity for various graph-related tasks. However, similar to deep neural networks, GNNs are also vulnerable to adversarial attacks. Empirical studies have shown that adversarially robust generalization has a pivotal role in establishing effective defense algorithms against adversarial attacks. In this paper, we contribute by providing adversarially robust generalization bounds for two kinds of popular GNNs, graph convolutional network (GCN) and message passing graph neural network, using the PAC-Bayesian framework. Our result reveals that spectral norm of the diffusion matrix on the graph and spectral norm of the weights as well as the perturbation factor govern the robust generalization bounds of both models. Our bounds are nontrivial generalizations of the results developed in (Liao et al., 2020) from the standard setting to adversarial setting while avoiding exponential dependence of the maximum node degree. As corollaries, we derive better PAC-Bayesian generalization bounds for GCN in the standard setting, which improve the bounds in (Liao et al., 2020) by avoiding exponential dependence on the maximum node degree.
keywords:
adversarially robust generalization bounds , PAC-Bayesian analysis , graph neural networks
journal: Neural Networks
[…]
organization=Center for Data Science, Zhejiang University, city=Hangzhou, country=P.R. China
@@ 1 Introduction
Graph neural network (GNN) (Scarselli et al., 2009) has developed rapidly in recent years, due to the excellent performance in a variety of applications with the graph-structure data. Generally, GNNs are neural networks that combine local aggregation of node features and non-linear transformations to predict the node-level labels or graph-level labels, in semi-supervised learning (Kipf and Welling, 2017) and supervised learning (Pan et al., 2015, 2016; Zhang et al., 2018; Ying et al., 2018), respectively. With the rapid development of GNN, there is an increasing interest in understanding the generalization (Scarselli et al., 2018; Verma and Zhang, 2019; Du et al., 2019; Garg et al., 2020; Liao et al., 2020; Esser et al., 2021; Ju et al., 2023) and other theoretical properties of GNN.
Szegedy et al. (2013) and Goodfellow et al. (2015) reveal that deep neural network (DNN) can be easily fooled by adversarial input samples and that adversarial training can improve the generalization of DNN. GNN also has a similar neural network structure and can be vulnerable to adversarial attacks on graph data. Previous work focuses on how to attack on GNN (Dai et al., 2018; Zügner et al., 2018; Bojchevski and Günnemann, 2019; Zügner et al., 2020; Liu et al., 2019) and how to defend GNN against adversarial attacks (Zhang and Zitnik, 2020; Wu et al., 2019; Entezari et al., 2020; Zhu et al., 2019; Tang et al., 2020), but the theoretical analysis is relatively limited.
A recent line of work discusses the generalization of GNN in the standard (non-adversarial) setting, and derive certain bounds on complexity notations including Vapnik–Chervonenkis (VC) dimension and Rademacher complexity that are used to characterize the generalization error, i.e., the disparity between the performance on training and test datasets of a model. Specifically, Scarselli et al. (2018) derive an upper bound of order $\mathcal{O}(p^{4}n^{2})$ for the VC dimension of GNNs with sigmoid activations, where $p$ is the number of parameters and $n$ is the number of nodes. Verma and Zhang (2019) derive a generalization bound for single-layer graph convolutional neural network (GCN) by analyzing the algorithmic stability of stochastic gradient descent, where the bound depends on the number of algorithm iterations and the largest absolute eigenvalue of the graph convolution filter. Du et al. (2019) consider an infinitely wide multi-layer GNN trained by gradient descent based on neural tangent kernels (Jacot et al., 2018), and derive a generalization bound for this particular model. These two results are constrained to specific settings and do not apply to other scenarios. Garg et al. (2020) present the pioneering data-dependent generalization bounds for GNNs employing Rademacher complexity, with a scaling factor of $\mathcal{O}\left(hd^{l}\sqrt{l\log\left(hd\right)}\right)$. Here, $d$ is the maximum degree of the graph, while $h$ and $l$ denote the width and depth of the network, respectively.
@@
In seeking to grasp the constrained robust generalization abilities of GNNs in the adversarial setting, in this paper, we provide adversarially robust generalization bounds for two kinds of popular GNNs, GCN (Kipf and Welling, 2017) and MPGNN (Dai et al., 2016; Jin et al., 2019), using PAC-Bayesian framework. Our result reveals that spectral norm of the diffusion matrix on the graph, spectral and Frobenius norm of the weights, as well as the perturbation factor govern the robust generalization bounds of both models. Our bounds for GCN complement the findings in the standard setting for DNN in (Neyshabur et al., 2017), as well as the results for GCN in (Liao et al., 2020). All the bounds of these studies have the same spectrally-normalized term of the weights and the dependencies on both depth and width:
@@
To our knowledge, our derived PAC-Bayesian adversarially robust generalization bounds may be the first kind for GNN in the adversarial robustness settings. Even in the standard setting, our bounds improve the bounds in (Liao et al., 2020) from $\mathcal{O}\left(ld^{\frac{l-1}{2}}\sqrt{h\log(lh)}\right)$ to $\mathcal{O}(l\sqrt{h\log(lh)})$ for GCN, and $\mathcal{O}\left(ld^{l-2}\sqrt{h\log(lh)}\right)$ to $\mathcal{O}(l\left\|P_{G}\right\|_{2}^{l-2}\sqrt{h\log(lh)})$ for MPGNN, avoiding the exponential dependence on the maximum degree $d$ of the graph. Meanwhile, our results remove the extra first-order and second-order Lipschitz-continuous assumptions in (Ju et al., 2023) for MPGNN, exhibiting comparable tightness on the dependency on the depth, width and maximum degree. In the context of adversarial robustness, the derived generalization bounds incorporate an additional factor $\epsilon$ that represents the intensity of the perturbation and is in general unavoidable. The derived bounds in the robust setting indicate that one could achieve roughly the same generalization performance as in the non-adversarial attack setting, provided that the $l_{2}$-norm attack level on the graph node, denoted as $\epsilon$, is less than the norm on the node. Additionally, according to our theory, the generalization error bounds do not have a significant effect on the graph edge attack. Therefore, one could overlook the attack on the graph edge and instead focus on the optimization error caused by the graph edge attack.
@@
The rest of the paper are organized as follows. In Section 2, we briefly review some other related works. In Section 3, we present the problem setup and preliminaries. We outline the main results in Section 4. Finally, we discuss the limitations and some open problems in Section 5.
@@ 2 Related Work
Generalization Bounds. Classical results rely on complexity notions of neural networks to regulate generalization bounds, employing measures such as VC-dimension as seen in (Anthony and Bartlett, 2009), or Rademacher complexity as discussed in (Bartlett and Mendelson, 2002). Norm-margin-based generalization bounds are developed using Rademacher complexity for fully-connected neural network in (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018). Neyshabur et al. (2015) establish a generalization bound with an exponential dependence $2^{l}$ on the network depth $l$ using Frobenius norms of parameters. Bartlett et al. (2017) showed a margin based generalization bound that depends on spectral norm and $1$-norm of the layers of the networks, reducing the dependence on depth. Golowich et al. (2018) provide bounds on their Rademacher complexity assuming norm constraints on the parameter matrix of each layer, refining the dependency on depth to $\sqrt{l}$. Neyshabur et al. (2017) establish a PAC-Bayesian framework to bound the generalization gap for fully-connected neural network, where the bound depends on the product of the spectral norm of the weights in each layer as well as the Frobenius norm of the weights. As mentioned above, the generalization bounds on GNN are derived using different tools such as VC-dimension (Scarselli et al., 2018), stability analysis (Verma and Zhang, 2019), Rademacher complexity (Garg et al., 2020) and PAC-Bayesian analysis (Liao et al., 2020; Ju et al., 2023).
@@
Adversarial Attack and Defense. Zügner et al. (2018) focus on the adversarial attacks against GCN and provide an algorithm to generate adversarial perturbations targeting on the node features and the graph structures. Dai et al. (2018) propose a reinforcement learning based attack method for adversarial attacks against GNN by modifying the structures of the graph. Zügner et al. (2020) develop a general attack on the training procedure of GCN using meta-gradients. Bojchevski and Günnemann (2019) propose a principled strategy for adversarial attacks on unsupervised node embeddings. There are also some researches focusing on the defense methods. Wu et al. (2019) provide a defense method by pre-processing the adjacency matrix of the graph. The methods in (Zhu et al., 2019; Entezari et al., 2020) defend against attacks by absorbing or dropping the noise information. Tang et al. (2020) use transfer learning to improve the robustness of GNN. Zhang and Zitnik (2020) develop a general algorithm to defend against a variety of training-time attacks that perturb the graph structure.
@@
Adversarial Generalization in DNN. A line of work presents the adversarially robust generalization bounds on DNNs using Rademacher complexity (Khim and Loh, 2018; Yin et al., 2019; Awasthi et al., 2020; Gao and Wang, 2021; Xiao et al., 2022). Khim and Loh (2018) derive robust generalization bounds for linear and neural network via Rademacher complexity of a transformed function. Yin et al. (2019) established generalization bounds for both linear models and one-hidden-layer neural networks based on the surrogate loss introduced in (Raghunathan et al., 2018) considering vector-$\infty$-norm-additive perturbation. Awasthi et al. (2020) give upper and lower bounds for Rademacher complexity of linear hypotheses with vector-$r$-norm-additive perturbation for $r\geq 1$, and also provide upper bounds on Rademacher complexity of one hidden layer neural networks, directly applying to the original network. Gao and Wang (2021) and Farnia et al. (2019) show robust generalization bounds for DNNs, considering specific attacks. Xiao et al. (2022) and Mustafa et al. (2022) employ two different techniques, respectively, to calculate the covering number that can be used to estimate the Rademacher complexity, for adversarial function classes of DNNs. Further, Xiao et al. (2023) present a robust generalization bound on DNN using PAC-Bayesian framework.
3 Preliminaries
@@
In this section, we introduce the problem setup for multi-class graph classification problem with GCN and MPGNN, and preliminary results on PAC-Bayesian analysis.
3.1 Problem Setup
@@
3.2 Background of PAC-Bayesian Analysis
@@
Neyshabur et al. (2017) provide the lemma to give a margin-based generalization bound derived from the PAC-Bayesian bound. Liao et al. (2020) improve the lemma and present the lemma in a two-side form. Both of the above lemmas are proposed in the standard settings. We present Lemma 3.2 of the two-side form with the robust margin loss in the adversarial settings. We provide proofs of Theorems 3.1 and 3.2 in Sections A.2.1 and A.2.2, respectively.
3.3 Assumptions
@@
4 Robust Generalization Bounds
In this section, we present our main results: improved generalization bounds in standard settings and adversarially robust generalization bounds for GCN and MPGNN.
4.1 Generalization Bounds for GCN
We first present the results for GCN in standard settings.
@@ 4.1 Generalization Bounds for GCN
The main difficulty is to well estimate the change in the robust margin operator with a perturbation on the parameters, adopting the spectral norm of $P_{G}$ rather than the maximum degree $d$. The estimating can be divided into two steps. The first step is to bound the change in the output of the model with the perturbation. We present the following lemma which shows that the change can be bounded by the norm of the parameters without dependence on the degree of the graph.
Lemma 4.3.
Let $f_{w}\in\mathcal{H}$ be a GCN model with $l$ layers. For any $w$, any perturbation $\Delta w=vec\left(\left\{\Delta W_{i}\right\}_{i=1}^{l}\right)$ such that for all $i\in[l]$, $\left\|\Delta W_{i}\right\|_{2}\leq\left\|W_{i}\right\|_{2}/l$, for any $G=(X,A)\in\mathcal{X}\times\mathcal{G}$, the change in the output of GCN is bounded as
Proof.
We denote the node representation in $j$-th ($j\leq l$) layer as
$f_{w}^{j}(G)=H_{j}=\sigma_{j}\left(P_{G}H_{j-1}W_{j}\right),\ j\in[l-1],$
$f_{w}^{l}(G)=H_{l}=\frac{1}{n}\bm{1}_{n}H_{l-1}W_{l}.$
Adding perturbation $\Delta w$ to the parameter $w$, that is, for the $j$-th ($j\leq l$) layer, the perturbed parameters are $W_{j}+\Delta W_{j}$ and denote $H_{j}^{\prime}=f_{w+\Delta w}^{j}(G),j\in[l]$.
Upper Bound on the Node Representation. For any $j<l$,
$\leq\left\|P_{G}H_{j-1}W_{j}\right\|_{F}$
$\leq\left\|P_{G}H_{j-1}\right\|_{F}\left\|W_{j}\right\|_{2}$
Upper Bound on the Change of Node Representation. For any $j<l$,
With Equation 3, we have
where the last inequality holds since $\left\|W_{i}+\Delta W_{i}\right\|_{2}\leq\left\|W_{i}\right\|_{2}+\left\|\Delta W_{i}\right\|_{2}\leq\left(1+1/l\right)\left\|W_{i}\right\|_{2}$ by the assumption that $\left\|\Delta W_{i}\right\|_{2}\leq\left\|W_{i}\right\|_{2}/l$.
Final Bound on the Readout Layer.
Using Equations 3 and 5, we have
$\leq\frac{1}{\sqrt{n}}\left\|X\right\|_{F}\left\|P_{G}\right\|_{2}^{l-1}\left[\left\|\Delta W_{l}\right\|_{2}\prod_{i=1}^{l-1}\left\|W_{i}\right\|_{2}\right.$
$\quad\left.+\left\|W_{l}+\Delta W_{l}\right\|_{2}\left(\prod_{i=1}^{l-1}\left\|W_{i}\right\|_{2}\right)\sum_{k=1}^{l-1}\frac{\left\|\Delta W_{k}\right\|_{2}}{\left\|W_{k}\right\|_{2}}\left(1+\frac{1}{l}\right)^{l-k-1}\right]$
$=\frac{1}{\sqrt{n}}\left\|X\right\|_{F}\left\|P_{G}\right\|_{2}^{l-1}\left(\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}\right)\left[\frac{\left\|\Delta W_{l}\right\|_{2}}{\left\|W_{l}\right\|_{2}}\right.$
$\quad\left.+\frac{\left\|W_{l}+\Delta W_{l}\right\|_{2}}{\left\|W_{l}\right\|_{2}}\sum_{k=1}^{l-1}\frac{\left\|\Delta W_{k}\right\|_{2}}{\left\|W_{k}\right\|_{2}}\left(1+\frac{1}{l}\right)^{l-k-1}\right]$
$\leq\frac{1}{\sqrt{n}}\left\|X\right\|_{F}\left\|P_{G}\right\|_{2}^{l-1}\left(\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}\right)\left[\frac{\left\|\Delta W_{l}\right\|_{2}}{\left\|W_{l}\right\|_{2}}\right.$
$\quad\left.+\left(1+\frac{1}{l}\right)\sum_{k=1}^{l-1}\frac{\left\|\Delta W_{k}\right\|_{2}}{\left\|W_{k}\right\|_{2}}\left(1+\frac{1}{l}\right)^{l-k-1}\right]$
$=\frac{1}{\sqrt{n}}\left\|X\right\|_{F}\left\|P_{G}\right\|_{2}^{l-1}\left(\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}\right)\left(1+\frac{1}{l}\right)^{l}\left[\frac{\left\|\Delta W_{l}\right\|_{2}}{\left\|W_{l}\right\|_{2}}\left(1+\frac{1}{l}\right)^{-l}\right.$
$\quad\left.+\sum_{k=1}^{l-1}\frac{\left\|\Delta W_{k}\right\|_{2}}{\left\|W_{k}\right\|_{2}}\left(1+\frac{1}{l}\right)^{-k}\right]$
where the last inequality holds since $1\leq\left(1+1/l\right)^{l}\leq e$. ∎
Lemma 4.4.
Under Assumptions 3.3 and 3.4, for any $w$ and any perturbation $\Delta w=vec\left(\left\{\Delta W_{i}\right\}_{i=1}^{l}\right)$ such that for all $i\in[l]$, $\left\|\Delta W_{i}\right\|_{2}\leq\left\|W_{i}\right\|_{2}/l$, the change in the margin operator is bounded as
Proof.
For any $i,j\in[K]$, by the definition of margin operator, we have
By Propositions A.2 and 4.3, we have
Further, let
$\quad\,\left|RM\left(f_{\tilde{w}}\left(G\right),i,j\right)-RM\left(f_{w}\left(G\right),i,j\right)\right|$
$=\left|\inf_{G^{\prime}\in\delta_{\tilde{w}}(G)}M\left(f_{\tilde{w}}\left(G^{\prime}\right),i,j\right)-\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G^{\prime}\right),i,j\right)\right|$
where the first inequality holds since, by the definition of $G(w)$,
Finally, the key to prove Theorem 4.2 using Lemmas 3.2 and 4.4, is to construct suitable prior and posterior distributions satisfying the conditions in the lemmas.
Proof of Theorem 4.2.
Let $\beta=\left(\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}\right)^{1/l}$. We normalize the weights as $\tilde{W}_{i}=\beta W_{i}/\left\|W_{i}\right\|_{2}$. Due to the homogeneity of ReLU, we have $f_{w}=f_{\tilde{w}}$. We can also verify that $\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}=\prod_{i=1}^{l}\left\|\tilde{W}_{i}\right\|_{2}$ and $\left\|W_{i}\right\|_{F}/\left\|W_{i}\right\|_{2}=\left\|\tilde{W}_{i}\right\|_{F}/\left\|\tilde{W}_{i}\right\|_{2}$, i.e., the terms appear in the bound stay the same after applying the normalization. Therefore, without loss of generality, we assume that the norm is equal across layers, i.e., $\forall i,\left\|W_{i}\right\|_{2}=\beta$.
Consider the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$ and the random perturbation $\Delta w\sim\mathcal{N}(0,\sigma^{2}I)$. Note that the $\sigma$ of the prior and the perturbation distributions are the same and will be set according to $\beta$. More precisely, we will set the $\sigma$ based on some approximation $\tilde{\beta}$ of $\beta$ since the prior distribution $P$ cannot depend on any learned weights directly. The approximation $\tilde{\beta}$ is chosen from a covering set which covers the meaningful range of $\beta$. For now, let us assume that we have a fixed $\tilde{\beta}$ and consider $\beta$ which satisfies $|\beta-\tilde{\beta}|\leq\beta/l$. Note that
where the last two inequalities hold since $\left(1-1/l\right)^{l-1}\geq 1/e$ and $\left(1+1/l\right)^{l}\leq e$. Indeed, using inequalities $\ln(x)\leq x-1,x>0$ and $\ln(x)\geq 1-1/x,x>0$, we can show that $l\ln\left(1+1/l\right)\leq 1$ and $(l-1)\ln\left(1-1/l\right)\geq(l-1)\left(1-\frac{1}{1-\frac{1}{l}}\right)=-1$.
From (Tropp, 2012), for $\Delta W_{i}\in\mathbb{R}^{h_{i-1}\times h_{i}}$ and $\Delta W_{i}\sim\mathcal{N}(0,\sigma^{2}I)$, we have
Taking a union bound, we have
$\geq 1-2lh\exp\left(-\frac{t^{2}}{2h\sigma^{2}}\right).$
Plugging this bound into Lemma 4.4, we have that with probability at least $1/2$ over $\Delta w$,
$\leq 2e(B+\epsilon)\beta^{l-1}l\sigma\sqrt{2h\log(4lh)}$
where we set $\sigma=\frac{\gamma}{4e^{2}(B+\epsilon)\tilde{\beta}^{l-1}l\sqrt{2h\log(4lh)}}$ to get the last inequality. Note that Lemma 4.4 also requires $\left\|\Delta W_{i}\right\|_{2}\leq\left\|W_{i}\right\|_{2}/l,\forall i\in[l]$. The requirement is satisfied if $\sigma\leq\frac{\beta}{l\sqrt{2h\log(4lh)}}$, which in turn can be satisfied if
Note that $\frac{\gamma}{4e(B+\epsilon)}\leq\beta^{l}$ is a sufficient condition of Equation 6. We will see how to satisfy this condition later.
We now compute the KL term in the PAC-Bayes bound in Lemma 3.2. As defined above, for a fixed $w$, the prior distribution $P=\mathcal{N}(0,\sigma^{2}I)$, the posterior distribution $Q$ is the distribution of $\Delta w$ shifted by $w$, that is, $Q=\mathcal{N}(w,\sigma^{2}I)$. Then we have,
From Lemma 3.2, fixing any $\tilde{\beta}$, with probability $1-\delta$ over $S$, for all $w$ such that $|\beta-\tilde{\beta}|\leq\beta/l$, we have,
since otherwise the bound holds trivially as $R_{\mathcal{D},0}\left(f_{w}\right)\leq 1$ by definition. If $\beta^{l}<\frac{\gamma}{2(B+\epsilon)}$, then for any $G$ and any $i,j\in[K]$, with Equation 3,
$\leq\frac{2}{\sqrt{n}}\left\|H_{l-1}^{*}\right\|_{F}\left\|W_{l}\right\|_{2}\leq 2(B+\epsilon)\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}$
where $\hat{G}=\arg\inf_{G^{\prime}\in\delta_{w}(G)}M\left(f_{w}\left(G^{\prime}\right),i,j\right)$ and $H_{l-1}^{*}$ is the corresponding node representation. Therefore, by the definition, we always have $\hat{R}_{S,\gamma}\left(f_{w}\right)=1$ when $\beta<\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}$.
Alternatively, if $\beta^{l}>\frac{\gamma\sqrt{m}}{2(B+\epsilon)}$, the term inside the big-O notation in Equation 7 would be
$\geq\sqrt{\frac{l^{2}h\log(lh)}{4}}\geq 1,$
where we use the facts that $\left\|W_{i}\right\|_{F}\geq\left\|W_{i}\right\|_{2}$ and we typically choose $h\geq 2$ and $l\geq 2$ in practice. Note that the lower bound in Equation 8 ensures that Equation 6 holds which in turn justifies the applicability of Lemma 4.4.
Consider a covering set centered at $C=\left\{\tilde{\beta}_{1},\dots,\tilde{\beta}_{|C|}\right\}$ of the interval in Equation 8 with radius $\frac{1}{l}\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}$, where $|C|$ is the size of the set. To make sure that $|\beta-\tilde{\beta}|\leq\left(\frac{\gamma}{2(B+\epsilon)}\right)^{1/l}/l\leq\beta/l$, the minimal size of $C$ should be
For a $\tilde{\beta}_{i}\in C$, we denote the event that Equation 7 holds for $|\beta-\tilde{\beta}|\leq\beta/l$ as $E_{i}$. Therefore, the probability of the event that Equation 7 holds for any $\beta$ satisfying Equation 8 is
$=\hat{L}_{S,\gamma}\left(f_{w}\right)+\mathcal{O}\left(\sqrt{\frac{(B+\epsilon)^{2}l^{2}h\log(lh)\prod_{i=1}^{l}\left\|W_{i}\right\|_{2}^{2}\sum_{i=1}^{l}\frac{\left\|W_{i}\right\|_{F}^{2}}{\left\|W_{i}\right\|_{2}^{2}}+\log(ml/\delta^{\prime})}{\gamma^{2}m}}\right).$
4.2 Generalization Bounds for MPGNN